Alvin is now 4 times as old as benjamin. Five years ago, the sum of their ages was 45 years. How old is each now?
Answer by addingup(3677)

A = 4B
A-5 + B-5 = 45
A + B = 55
4B + B = 55
5B = 55
B = 11
Benjamin is 11 and Alvin is 4 times as old, 11 x 4 = 44